OpenStudy (anonymous):

The distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 306 miles in 8.5 hours, how many miles will Olivia travel in 9.1 hours at the same constant speed?

OpenStudy (anonymous):

306/8.5 = x/9.1 (306 miles to 8.5 hrs = x miles to 9.1 hrs) Now cross multiply (8.5)(x) = (306)(9.1) solve for x

OpenStudy (anonymous):

D = vt D1 = 306 miles , t1 = 8.5 hours constant ,v = 306 / 8.5 NOW if t2 = 9.1 hours , D2 = v t = ( 306/8.5) x 9.1 = ____

OpenStudy (anonymous):

Either way will give you the same answer

OpenStudy (anonymous):

Ok thank you:)

OpenStudy (anonymous):

no problem :)